# CAST: Contrastive Adaptation and Distillation for Semi-Supervised Instance Segmentation

### v3

@@ 4.4 Ablation Studies
We perform a series of ablation experiments on the Cityscapes dataset to isolate the contributions of different components in the CAST pipeline and the pixel-level contrastive loss.
Impact of Loss Components. In the knowledge distillation stage, the overall loss is a weighted sum of three components: the supervised loss ($\mathcal{L}_{\text{sup}}$), the semi-supervised pseudo-label loss ($\mathcal{L}_{\text{semi}}$), and the pixel-level self-supervised contrastive loss ($\mathcal{L}_{\text{pxl}}$). Table 2 presents an ablation study assessing the contribution of each loss term. The results indicate that adding the pseudo-label loss substantially improves performance over the supervised baseline, and further inclusion of the self-supervised loss leads to the highest overall performance.
Table 2: Loss Ablation: Impact of supervised ($\mathcal{L}_{\rm sup}$), pseudo-label ($\mathcal{L}_{\rm semi}$), and pixel-level ($\mathcal{L}_{\rm pxl}$) terms on teacher vs. student maskAP.
Method $\mathcal{L}_{\rm sup}$ $\mathcal{L}_{\rm semi}$ $\mathcal{L}_{\rm pxl}$ maskAP (%)
Teacher Student
(a) Supervised Only 28.7 21.1
(b) + Pseudo-Label 29.7 30.7
(c) + pixel-level loss 29.6 27.5
(d) + (b) + (c) 30.5 32.2
Ablation of Negative Sampling via Various Probability Maps.
To validate our negative sampling strategy in the pixel-level contrastive loss, Table 3 compares four sampling methods: Uniform: negatives sampled uniformly across the image; Mask-Only: The probability map is derived solely from mask predictions, with class probabilities assumed to be uniform. Class-Only: The map is generated only from class predictions, assuming a uniform spatial distribution for the mask. Fusion: Combining both mask and class predictions. The fusion strategy achieves the best results, with 32.2 maskAP and 56.5 AP50.
(a) Negative Sampling Strategies
Method maskAP (%) AP50 (%)
@@
Mask-Only $30.6$ $55.0$
Class-Only $31.1$ $55.3$
@@
(b) Schematic of Sampling Distributions
Table 3: Ablation of Negative Sampling Strategies on Cityscapes. (3) Quantitative results for uniform, mask-only, class-only, and fusion samplers (maskAP and AP50). (3) Schematic sketch of the corresponding pixel-level sampling probability distributions.
Hyperparameter Sensitivity. We evaluate CAST’s sensitivity to three key hyperparameters on Cityscapes: contrastive weight $\lambda_{\rm pxl}$, negatives per anchor $K$, and temperature $T$, by measuring both teacher and student mask AP (%) and AP50 (%). Table 4 reports the full sweep. We find that $\lambda_{\rm pxl}=0.2$ and $T=0.2$ consistently maximize performance. For the number of negatives, $K=256$ offers the best trade-off: although $K=512$ yields a slight increase in teacher maskAP (30.9 vs. 30.5) and AP50 (57.1 vs. 56.6), and comparable student metrics, the marginal gains saturate relative to the increased sampling cost. Therefore, we adopt $K=256$ throughout.
@@
Student Architecture Variants.
We evaluate two design axes for the student model under CAST distillation protocol: (i) the encoder backbone (with a fixed DPT decoder), and (ii) the decoder head (with a fixed DINOv2-S encoder). Table 5 reports accuracy along with parameter counts, on the Cityscapes validation set. The combination of DINOv2-S encoder and DPT head achieves the best accuracy with a compact footprint.
@@
Encoder maskAP AP50 Params (M)
@@
Decoder maskAP AP50 Params (M)
@@ Appendix 5 Conclusions
We have introduced CAST, a rigorously designed SSKD pipeline that fuses self-training, instance-aware pixel-wise contrastive learning, and final supervised finetuning to compress large VFMs into compact student experts with comparable performance. Empirically, our 11$\approx$ 11$\times$ 11 × smaller student exceeds its adapted teacher by +3.4 maskAP in Cityscapes and +1.5 maskAP in ADE20K, while cutting compute and parameter counts demonstrating that dense contrastive supervision can unlock substantial gains in low-label regimes. Our theoretical analysis further guarantees that our negative sampling scheme provably increases inter-instance margins under mild assumptions. Looking forward, streamlining CAST into a single unified objective, extending its evaluation to diverse domains, and integrating uncertainty quantification will be critical steps toward safe, equitable, and broadly deployable segmentation solutions.
@@
Broader Impacts. This paper focuses on developing high fidelity instance segmentation models under tight annotation and compute budgets, running efficiently on workstation GPUs and making it practical for domains such as autonomous driving, mobile and industrial robotics, augmented reality and agricultural or infrastructure monitoring. These safety critical applications are especially vulnerable to biases inherited from generalized VFMs (e.g., underrepresented object classes or ambiguous category definitions, which can lead to degraded performance or unfair outcomes. By distilling task-specific expert models, CAST helps reduce these generic failure modes, but careful domain specific validation, bias auditing, and uncertainty quantification remain essential for deployment. Limitations. While our CAST framework achieves significant performance improvements, the training pipeline comprising multiple stages introduces additional complexity. Future work could focus on streamlining these stages and further reducing the training cost.
@@
Figure 7: Qualitative bias reduction in stage-wise distillation. Top row: pseudo-labels generated by the pretrained teacher. Bottom row: student predictions after distillation and refinement, demonstrating reduced pseudo-label bias and sharper instance boundaries.
@@ Appendix 6 Dataset Splits
Table 6 summarizes the datasets used in our experiments. We use a 10% labelled split of Cityscapes’ 2 975 training images (298 labeled / 2 677 unlabeled) and a stratified 20% split of ADE20K’s 20 210 training images (1 000 labeled / 2 537 unlabeled). Standard validation sets are retained (500 images for Cityscapes, 2 000 for ADE20K). Exact image-ID lists will be released with our code.
@@ Appendix 7 Hyperparameters
Let $z_{a}$, $z^{+}$ and $\{z^{-}_{r}\}_{r=1}^{R}$ be the unit norm embeddings of an anchor pixel, its positive, and $R$ negatives. Define
@@
By Assumption 3.1, each negative embedding $z^{-}_{r}$ is inter-instance with probability $p$, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 0$, and intra-instance with probability $1-p$, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 1$. Hence
@@ Appendix 10 Additional Ablations
Replacing our asymmetric InfoNCE (§3.2) with an margin-based hinge loss (margin = 0.2) yields identical maskAP (32.2%) and +0.6 AP50, at the cost of $1.6$× longer training. This evaluates whether enforcing a fixed positive–negative margin can match or improve upon the performance of InfoNCE.
@@ 10.2 Ablation: Debias Score Formulation
[…]
Table 9: Debias Score Formulation Ablation. Original vs. squared vs. square-root debias scores.
@@ 10.3 Ablation: Negative Sampling Scope
Mini-batch only 32.2 56.5
@@
Sampling from a memory bank of 10 k embeddings yields a modest performance gain (+0.5 maskAP, +0.8 AP50) compared to in-batch sampling. However, this approach incurs approximately $2.2$× longer training time due to the overhead of maintaining and querying the memory bank.

### v4

@@ 4.4 Ablation Studies
We perform a series of ablation experiments to isolate the contributions of each component in the CAST pipeline. These include analyses of loss functions, training stages, negative sampling strategies, hyperparameters, and student architecture choices.
Impact of Loss Components. During distillation, the objective combines three terms: supervised loss ($\mathcal{L}_{\text{sup}}$), semi-supervised pseudo-label loss ($\mathcal{L}_{\text{semi}}$), and pixel-level self-supervised contrastive loss ($\mathcal{L}_{\text{pxl}}$). Table 2 shows that adding $\mathcal{L}_{\text{semi}}$ improves student performance from 21.1 to 30.7 maskAP, while further including $\mathcal{L}_{\text{pxl}}$ yields the best result of 32.2 maskAP, confirming complementary benefit.
Table 2: Ablations on Cityscapes (10% labels). Left: effect of loss terms. Right: effect of CAST stages.
Method $\mathcal{L}_{\rm sup}$ $\mathcal{L}_{\rm semi}$ $\mathcal{L}_{\rm pxl}$ Teacher Student
(a) Sup. only 28.7 21.1
(b) + Pseudo 29.7 30.7
(c) + Pixel loss 29.6 27.5
(d) (b)+(c) 30.5 32.2
(a) Loss ablation
Variant Teacher Adapt. Distill. Student FT maskAP
Full CAST 33.9
No Student FT 32.2
No Teacher Adapt. 25.7
Distillation Only 23.8
No Distill. (Sup.) 21.1
(b) Stage ablation
Impact of Training Stages. Beyond the contribution of individual loss terms, we further ablate each stage of CAST to justify their necessity. Table 2 shows results on Cityscapes (10% labels), where we drop exactly one stage at a time.
The supervised baseline achieves 21.1 maskAP. Adding distillation alone improves this to 23.8 (+2.7), and further adding student fine-tuning raises it to 32.2 (+8.4). Without teacher adaptation, performance drops to 25.7, underscoring the need to align the teacher with the target domain. The full three-stage CAST pipeline achieves best result of 33.9 maskAP, a +12.8 improvement over baseline.
Ablation of Negative Sampling via Various Probability Maps. To validate our negative sampling strategy in the pixel-level contrastive loss, Table 3 compares four sampling methods: Uniform: negatives sampled uniformly across the image; Mask-Only: The probability map is derived solely from mask predictions, with class probabilities assumed to be uniform. Class-Only: The map is generated only from class predictions, assuming a uniform spatial distribution for the mask. Fusion: Combining both mask and class predictions. The fusion strategy achieves the best results, with 32.2 maskAP and 56.5 AP50.
(c) Negative Sampling Strategies
Method maskAP (%) maskAP50 (%)
@@
Mask‐Only $30.6$ $55.0$
Class‐Only $31.1$ $55.3$
@@
(d) Schematic of Sampling Distributions
Table 3: Ablation of Negative Sampling Strategies on Cityscapes. (3) Quantitative results for uniform, mask‐only, class‐only, and fusion samplers (maskAP and maskAP50). (3) Schematic sketch of the corresponding pixel‐level sampling probability distributions.
Hyperparameter Sensitivity. We evaluate CAST’s sensitivity to three key hyperparameters on Cityscapes: contrastive weight $\lambda_{\rm pxl}$, negatives per anchor $K$, and temperature $T$, by measuring both teacher and student maskAP (%) and maskAP50 (%). Table 4 reports the full sweep. We find that $\lambda_{\rm pxl}=0.2$ and $T=0.2$ consistently maximize performance. For the number of negatives, $K=256$ offers the best trade-off: although $K=512$ yields a slight increase in teacher maskAP (30.9 vs. 30.5) and maskAP50 (57.1 vs. 56.6), and comparable student metrics, the marginal gains saturate relative to the increased sampling cost. Therefore, we adopt $K=256$ throughout.
@@
Student Architecture Variants. We evaluate two design axes for the student model under CAST distillation protocol: (i) the encoder backbone (with a fixed DPT decoder), and (ii) the decoder head (with a fixed DINOv2-S encoder). Table 5 reports accuracy along with parameter counts, on the Cityscapes validation set. The combination of DINOv2-S encoder and DPT head achieves the best accuracy with a compact footprint.
@@
Encoder maskAP maskAP50 Params (M)
@@
Decoder maskAP maskAP50 Params (M)
@@
Scalability with Labeled Fractions. We evaluate CAST under different fractions of labeled data to assess scalability in semi-supervised settings. Following the protocol in [3], we train with 5%, 10%, and 30% labeled splits of Cityscapes. As shown in Table 6, CAST consistently outperforms prior methods across all fractions. At 5% labels, CAST achieves 30.7 AP, far exceeding PAIS (18.0) and Guided Distillation (23.0). At 30% labels, CAST reaches 40.4 AP, surpassing the strongest baseline (37.8 from S4M) by +2.6 AP. These results demonstrate that CAST remains effective under scarce supervision while scaling gracefully with additional labeled data.
Table 6: Scalability across label fractions on Cityscapes. Results with 5%, 10%, and 30% labeled data.
Dataset Fraction Teacher Adapt. Distillation CAST (student) PAIS [17] Guided dist. [3] S4[44]
5% 29.4 29.2 30.7 18.0 23.0 30.1
10% 30.5 32.2 33.9 22.9 30.8 33.3
30% 33.3 38.5 40.4 32.8 35.6 37.8
Additional ablations, including teacher adaptation variants, loss formulations, sampling scope, and backbone comparisons, are provided in the supplementary material (Section E).
@@ Appendix 5 Conclusions
We have introduced CAST, a rigorously designed SSKD pipeline that fuses self‐training, instance-aware pixel‐wise contrastive learning, and final supervised finetuning to compress large VFMs into compact student experts with comparable performance. Empirically, our $\approx 11\times$ smaller student exceeds its adapted teacher by +3.4 maskAP in Cityscapes and +1.5 maskAP in ADE20K, while cutting compute and parameter counts demonstrating that dense contrastive supervision can unlock substantial gains in low‐label regimes. Our theoretical analysis further guarantees that our negative sampling scheme provably increases inter‐instance margins under mild assumptions. Looking forward, streamlining CAST into a single unified objective, extending its evaluation to diverse domains, and integrating uncertainty quantification will be critical steps toward safe, equitable, and broadly deployable segmentation solutions.
@@
Figure 7: Qualitative bias reduction in stage‐wise distillation. Top row: pseudo‐labels generated by the pretrained teacher. Bottom row: student predictions after distillation and refinement, demonstrating reduced pseudo‐label bias and sharper instance boundaries.
Ethics Statement
This work does not involve human subjects, private data, or sensitive content. All datasets used are publicly available. Portions of the manuscript were polished using large language model (LLM) for clarity; this use was limited to text editing and did not affect the research process, experiments, or results.
Reproducibility Statement
We provide detailed descriptions of datasets, model architectures, and training procedures to ensure reproducibility. All datasets (Cityscapes, ADE20K) are publicly available, and the 5%, 10%, and 30% labeled splits follow established protocols (Section 4). Implementation details, including hyperparameters, software environment, and GPU usage, are reported in Section 4.2, with extensive ablations and sensitivity analyses in Appendix E. To facilitate further research, we will release our code upon the completion of the anonymous review process.
@@ Appendix 6 Dataset Splits
Table 7 summarizes the datasets used in our experiments. We use a 10% labelled split of Cityscapes’ 2 975 training images (298 labeled / 2 677 unlabeled) and a stratified 20% split of ADE20K’s 20 210 training images (1 000 labeled / 2 537 unlabeled). Standard validation sets are retained (500 images for Cityscapes, 2 000 for ADE20K). Exact image‐ID lists will be released with our code.
@@ Appendix 7 Hyperparameters
Let $_{a}, z+z^{+}$ and $\{z^{-}_{r}\}_{r=1}^{R}$ be the unit norm embeddings of an anchor pixel, its positive, and $R$ negatives. Define
@@
By Assumption 3.1, each negative embedding $z^{-}_{r}$ is inter-instance with probability $p$, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 0$, and intra-instance with probability 1p1-p, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 1$. Hence
@@ Appendix 10 Additional Ablations
Replacing our asymmetric InfoNCE (§3.2) with an margin-based hinge loss (margin = 0.2) yields identical maskAP (32.2%) and +0.6 AP50, at the cost of 1.61.6× longer training. This evaluates whether enforcing a fixed positive–negative margin can match or improve upon the performance of InfoNCE.
@@ 10.2 Ablation: Debias Score Formulation
We evaluate three instantiations of the debias score function $s^{\mathit{deb}}$ (§3.2):
  • Original $s^{\mathit{deb}}$: fusion of mask and class confidences (ours).
  • $\bigl(s^{\mathit{deb}}\bigr)^{2}$: square each score to amplify high‐confidence negatives.
  • $\sqrt{s^{\mathit{deb}}}$: take the square root of each score to temper the bias.
Table 10: Debias Score Formulation Ablation. Original vs. squared vs. square‐root debias scores.
@@ 10.3 Ablation: Negative Sampling Scope
Mini‐batch only 32.2 56.5
@@
Sampling from a memory bank of 10 k embeddings yields a modest performance gain (+0.5 maskAP, +0.8 AP50) compared to in‐batch sampling. However, this approach incurs approximately 2.22.2× longer training time due to the overhead of maintaining and querying the memory bank.